OpenStudy (anonymous):

Use an inequality to solve the problem. Be sure to show the inequality and all of your work. One side of a rectangle is 8 inches and the other side is x inches. What values of x will make the perimeter at least 46 inches?

OpenStudy (dumbcow):

perimeter is the sum of all the sides P = 8+8+x+x = 16+2x Now set that greater than or equal to 46 --> 16+2x >= 46 solve for x